### v1

@@ 5.1.2 Experimental Results
Troja et al [41] 1-DB Peer-to-Peer $1650$ ms $11760$ ms $13410$ ms $12$ MB
Li et al [29] 1-DB Pseudo-ID WiFi AP+Loc.Server $210$ ms
Xin et al [47] 1-DB PIR WiFi AP+QRA $292.8$ ms $142.7$ ms $407.4$ ms $430.1$ ms 325KB
LP-Chor [23] $\ell$-DB PIR $7.7$ ms $480$ ms $620$ ms $753$ KB
LP-Goldberg [23] $\ell$-DB PIR $320$ ms $1210$ ms $1780$ ms 66 MB
RAID-LP-Chor [23] $\ell$-DB PIR $0.4$ ms $22$ ms $210$ ms $125$ KB
Zeng et al [48] 1-DB BS+ECC PseudoID $87$ ms $27$ ms $135$ ms $1.24$ KB
TrustSAS [24] $\ell$-DB PIR EPID $329.4$ ms $324.6$ ms $4954$ ms $1.25$ MB
PACDoSQ [14] $\ell$-DB PIR Tor HBP $28.1$ ms $199$ ms $1373.6$ ms $605.92$ KB
@@
  • Libraries: Virtual Machines running Ubuntu simulated PIR costs, using the percy++ library [^8] https://percy.sourceforge.net/ for multi-server PIR, the Open Quantum-Safe library [^9] https://openquantumsafe.org/ for PQC primitives, and OpenSSL for cryptographic operations and arithmetic. Variables: We consider six databases for multi-DB schemes with $|DB|=560~\textit{MB}$ and $400$ rows/columns as described in [47]. Key terms include BS (base station), HBP (hash-based puzzles), QRA (quadratic residue assumption), and EPID (enhanced privacy ID based on direct anonymous attestation).
@@ 5.2 Network-Level Evaluation
This section presents an instantiation of the SLAPX protocol in a realistic wireless network setting and evaluates its performance through simulations. We use ns3 [1] to model the deployment of SLAPX and analyze its communication overhead, while the 5G NR (New Radio) LENA module [37] is employed to simulate the 5G infrastructure.
@@ 5.2.1 Simulation Setup and Configuration
5G Parameter Configuration: The 5G NR LENA module offers a wide range of configuration options for realistic 5G simulations. For simplicity, we retain default settings for most parameters and manually configure only a selected subset. The configured parameters and their values are summarized in Table 3. Specifically, the channel model follows the standard NR specification [2] used by default in the LENA module. The transmission power of both UE and gNodeB nodes is set to $30$ dBm, reflecting typical deployment conditions. The Maximum Transmission Unit (MTU) is set to the standard Ethernet frame size of $1500$ bytes, and the backhaul network is modeled with a data rate of $1$ Gbps and an end-to-end latency of $2$ ms.
@@
Server Modeling: Because ns3 is a discrete-event simulator, directly modeling real-time server overload caused by DoS attacks is nontrivial. While network-level congestion can be simulated, capturing the impact of DoS attacks on server-side processing requires an abstracted model. Therefore, each server is represented using a queue-and-worker abstraction, a common approach in analytical DoS modeling [44]. Specifically, the server consists of two components: (i) Workers, which model independent computational units (e.g., CPU cores) capable of processing requests in parallel; and (ii) Queue, which buffers incoming requests when all workers are busy. However, when the queue is full, incoming requests are dropped. This model enables us to capture server saturation and request loss under DoS conditions, serving as the basis for evaluating DoS scenarios in our simulations.
@@ 5.2.2 Simulation Results
Fragmentation Analysis of SLAPX: Based on a realistic 5G deployment, we evaluate SLAPX through numerical analysis focusing on fragmentation behavior and communication overhead. Due to the relatively large size of certain protocol messages, packet fragmentation becomes observable as illustrated in Figure 5, where the upper illustration shows the relationship between MTU values and fragmentation across different SLAPX message types. Although the 5G standard does not mandate a fixed MTU, many deployments adopt the standard Ethernet MTU of $1500$ bytes [18]. Under this setting, only the Service Request and Spectrum Request messages experience fragmentation, increasing from one to two packets per message. The lower plot in Figure 5 further analyzes communication overhead under varying MTU sizes, illustrating the proportion of protocol and network headers relative to payload size. As expected, messages with smaller payloads incur higher relative overhead due to header dominance.
@@
Distance Hijacking: For one of the location spoofing scenarios, we evaluate the effectiveness of the Proximity Verification function (ProxVerify), which enables a gNodeB to estimate the distance to UEs using a combination of received signal strength (RSS) and round-trip time (RTT) measurements. In the considered distance hijacking attack, a malicious UE communicates with a compromised benign UE over an ad-hoc WiFi network using CSMA, while the benign UE maintains a legitimate 5G NR connection with the gNodeB. The benign UE is configured to relay all messages between the malicious UE and the gNodeB without delay. The proximity threshold is set to 50 meters, in accordance with FCC guidelines [19], such that devices within this range are classified as proximate, while those beyond it are considered distant.
@@ 6 Related Work
(i) Location Privacy and Anonymity in DB-CRN: In DB-CRNs, regulatory requirements mandate that unlicensed users (i.e., SUs) disclose precise location information, device characteristics, and spectrum requests to geolocation databases in order to obtain authorized channels and transmission parameters [4]. This regulatory enforcement fundamentally conflicts with the principle of privacy, as users must sacrifice their privacy for spectrum access, thereby exposing mobility patterns and long-term behavioral profiles, which then enables tracking and profiling attacks [24]. Moreover, location privacy without anonymity or unlinkability remains insufficient, as stable identifiers combined with coarse location allow tracking and linkage across queries [33]. Standard protections (e.g., TLS) do not prevent legitimate entities from observing, storing, and correlating sensitive query metadata [39], either. Furthermore, the loss of anonymity expands the attack surface and can facilitate active threats such as selective DoS, targeted blocking, and user discrimination [45, 14].
@@
Despite extensive research efforts, existing privacy-enhancing mechanisms remain inadequate for real-world DB-CRN deployments. Schemes relying on k-anonymity or pseudonymous identifiers [49] provide only weak protection and lack rigorous security guarantees unless very large anonymity sets are maintained—an assumption that rarely holds in dense, mobile, or highly dynamic environments. Techniques based on location obfuscation or differential privacy [42], while theoretically grounded, intentionally perturb location data and consequently impair the precision and dependability of spectrum availability decisions. Cryptographic approaches leveraging Private Information Retrieval (PIR) [24, 47] offer stronger privacy guarantees but incur prohibitive computational and communication overhead, hindering scalability and practical deployment, while typically assuming benign user behavior and neglecting adversarial users. Consequently, our proposed solution utilizes primitives such as DAC and RLRS to provide effective protection in DB-CRNs through lightweight and scalable mechanisms that jointly ensure location privacy, anonymity, and unlinkability, while remaining compatible with regulatory mandates and operating in resource-constrained environments.
@@
(ii) Location Proof and Spoofing Resistance in DB-CRN: DB-CRNs fundamentally operate as location-based systems in which spectrum availability decisions depend on real-time, user-provided location information stored in geolocation DBs. Consequently, spoofed or falsified locations undermine the basis of regulatory authorization and can cause disproportionate harm, as attackers may impersonate legitimate entities or manipulate location and usage data to obtain unauthorized spectrum access, resulting in spectrum interference, service disruption, and economic loss [34].
@@
While some approaches introduce location-proof mechanisms, they often rely on strong trust assumptions such as trusted infrastructure, specialized hardware, or dedicated location-verification servers, which are impractical in rural, ad hoc, or infrastructure-limited deployments [47]. The need for location verification becomes even more critical when combined with location privacy and anonymity, creating a fundamental paradox in DB-CRNs: revealing precise location information to prove correctness undermines privacy, while hiding location without verification enables spoofing. Thus, many existing solutions compromise user privacy by exposing fine-grained location data to verifiers, enabling tracking and linkage. These shortcomings highlight the necessity for a practical, privacy-preserving location verification framework that can authenticate user proximity without disclosing sensitive location data, while remaining robust against diverse spoofing and distance-fraud attacks under realistic DB-CRN deployment conditions [15].
@@
(iii) DoS Countermeasures and Spectrum Management for Next-Gen Network Systems: The centralized geolocation databases in DB-CRNs, due to their frequent interaction with massive IoT deployments, result in high-value choke points whose unavailability can disrupt spectrum access at scale [26]. Due to the cost asymmetry, even weak attackers can overwhelm spectrum DBs with excessive or illegitimate queries, causing high verification and processing overhead for servers, which in turn degrades spectrum coordination and reduces service availability.
Although various countermeasures have been proposed, including intrusion detection systems (IDSs), blockchain-based access control, and cryptographic Client Puzzle Protocols (CPPs), most existing defenses remain reactive and focus primarily on attack detection rather than prevention [13]. Such approaches often require continuous monitoring of traffic patterns, which incurs substantial operational overhead, and remain vulnerable to adversarial evasion [17]. CPP-based defenses impose computational costs on clients prior to service access [9], thereby limiting the rate of malicious requests. However, at large scale, they introduce new bottlenecks related to puzzle generation, distribution, and parallelization, which can burden both servers and legitimate users [5]. While outsourcing puzzle generation can alleviate server-side overhead [14], it merely shifts the attack surface to spectrum databases, which remain vulnerable to query flooding.
@@ A.1 Security Proofs
(i) Client Anonymity and Untraceability: SLAPX ensures client anonymity and untraceability via credential unforgeability and credential anonymity (Definitions 5 and 6). For any PPT adversary $\mathcal{A}$ interacting with issuance, delegation, and presentation oracles, the advantage in distinguishing which of two honest users generated a valid presentation satisfies $Adv_{\mathcal{A}}^{\mathrm{anon}}(\lambda):=\left|\Pr[b^{\prime}=b]-\tfrac{1}{2}\right|\leq negl(\lambda)$, where $b$ is the challenger’s hidden bit selecting the presenting user. Otherwise, $\mathcal{A}$ can be reduced to either (i) breaking the knowledge soundness of the underlying ZKPoK, (ii) distinguishing re-randomized structure-preserving signatures, or (iii) distinguishing public-key–switched presentations, yielding a distinguisher for the Decisional Diffie–Hellman (DDH) problem in the underlying group $\mathbb{G}$. In particular, the SPSEQ-UC scheme guarantees origin-hiding, derivation-privacy, and conversion-privacy [20, 32], implying that randomized signatures, delegated commitment vectors, and key-switched credentials are computationally indistinguishable from fresh ones. These properties compose under repeated application, ensuring that credential presentations remain anonymous and unlinkable with all but negligible probability.
(ii) Client Location Privacy: SLAPX preserves user location privacy through credential unlinkability (Definition 7). In particular, location information is protected by the unlinkability of signature–commitment pairs generated via the SPSEQ-UC scheme [32]. The scheme supports signature re-randomization and public-key switching, ensuring that multiple presentations of the same underlying commitment yield computationally independent transcripts. Formally, for any PPT adversary $\mathcal{A}$ interacting with the proving oracle and observing a sequence of credential showings associated with the same location attribute $\ell$, the advantage of linking two presentations satisfies $Adv_{\mathcal{A}}^{\mathrm{link}}(\lambda):=\Big|\Pr[b^{\prime}=b]-\tfrac{1}{2}\Big|\leq negl(\lambda)$, where $b$ denotes the challenge bit selecting which of two unlinkable presentations is shown. Otherwise, $\mathcal{A}$ can be used to distinguish re-randomized SPSEQ-UC signatures or correlate switched public keys, contradicting the unlinkability of SPSEQ-UC under the generic group model. Consequently, repeated spectrum queries and location proofs in SLAPX remain computationally indistinguishable and unlinkable to verifiers, ensuring location privacy with all but negligible probability.

### v2

@@ 5.1.2 Experimental Results
Troja et al [troja2014leveraging] 1-DB Peer-to-Peer $1650$ ms $11760$ ms $13410$ ms $12$ MB
Li et al [li2015privacy] 1-DB Pseudo-ID WiFi AP+Loc.Server $210$ ms
Xin et al [xin2016privacy] 1-DB PIR WiFi AP+QRA $292.8$ ms $142.7$ ms $407.4$ ms $430.1$ ms 325KB
LP-Chor [grissa2019location] $\ell$-DB PIR $7.7$ ms $480$ ms $620$ ms $753$ KB
LP-Goldberg [grissa2019location] $\ell$-DB PIR $320$ ms $1210$ ms $1780$ ms 66 MB
RAID-LP-Chor [grissa2019location] $\ell$-DB PIR $0.4$ ms $22$ ms $210$ ms $125$ KB
Zeng et al [zeng2019efficient] 1-DB BS+ECC PseudoID $87$ ms $27$ ms $135$ ms $1.24$ KB
TrustSAS [grissa2021anonymous] $\ell$-DB PIR EPID $329.4$ ms $324.6$ ms $4954$ ms $1.25$ MB
PACDoSQ [darzi2024privacy] $\ell$-DB PIR Tor HBP $28.1$ ms $199$ ms $1373.6$ ms $605.92$ KB
@@
  • Libraries: Virtual Machines running Ubuntu simulated PIR costs, using the percy++ library [^8] https://percy.sourceforge.net/ for multi-server PIR, the Open Quantum-Safe library [^9] https://openquantumsafe.org/ for PQC primitives, and OpenSSL for cryptographic operations and arithmetic. Variables: We consider six databases for multi-DB schemes with $|DB|=560~\textit{MB}$ and $400$ rows/columns as described in [xin2016privacy]. Key terms include BS (base station), HBP (hash-based puzzles), QRA (quadratic residue assumption), and EPID (enhanced privacy ID based on direct anonymous attestation).
@@ 5.2 Network-Level Evaluation
This section presents an instantiation of the SLAPX protocol in a realistic wireless network setting and evaluates its performance through simulations. We use ns3 [NS3] to model the deployment of SLAPX and analyze its communication overhead, while the 5G NR (New Radio) LENA module [5G_LENA] is employed to simulate the 5G infrastructure.
@@ 5.2.1 Simulation Setup and Configuration
5G Parameter Configuration: The 5G NR LENA module offers a wide range of configuration options for realistic 5G simulations. For simplicity, we retain default settings for most parameters and manually configure only a selected subset. The configured parameters and their values are summarized in Table 3. Specifically, the channel model follows the standard NR specification [3GPPChannelModel] used by default in the LENA module. The transmission power of both UE and gNodeB nodes is set to $30$ dBm, reflecting typical deployment conditions. The Maximum Transmission Unit (MTU) is set to the standard Ethernet frame size of $1500$ bytes, and the backhaul network is modeled with a data rate of $1$ Gbps and an end-to-end latency of $2$ ms.
@@
Server Modeling: Because ns3 is a discrete-event simulator, directly modeling real-time server overload caused by DoS attacks is nontrivial. While network-level congestion can be simulated, capturing the impact of DoS attacks on server-side processing requires an abstracted model. Therefore, each server is represented using a queue-and-worker abstraction, a common approach in analytical DoS modeling [Wang2007]. Specifically, the server consists of two components: (i) Workers, which model independent computational units (e.g., CPU cores) capable of processing requests in parallel; and (ii) Queue, which buffers incoming requests when all workers are busy. However, when the queue is full, incoming requests are dropped. This model enables us to capture server saturation and request loss under DoS conditions, serving as the basis for evaluating DoS scenarios in our simulations.
@@ 5.2.2 Simulation Results
Fragmentation Analysis of SLAPX: Based on a realistic 5G deployment, we evaluate SLAPX through numerical analysis focusing on fragmentation behavior and communication overhead. Due to the relatively large size of certain protocol messages, packet fragmentation becomes observable as illustrated in Figure 5, where the upper illustration shows the relationship between MTU values and fragmentation across different SLAPX message types. Although the 5G standard does not mandate a fixed MTU, many deployments adopt the standard Ethernet MTU of $1500$ bytes [RFC_MTU]. Under this setting, only the Service Request and Spectrum Request messages experience fragmentation, increasing from one to two packets per message. The lower plot in Figure 5 further analyzes communication overhead under varying MTU sizes, illustrating the proportion of protocol and network headers relative to payload size. As expected, messages with smaller payloads incur higher relative overhead due to header dominance.
@@
Distance Hijacking: For one of the location spoofing scenarios, we evaluate the effectiveness of the Proximity Verification function (ProxVerify), which enables a gNodeB to estimate the distance to UEs using a combination of received signal strength (RSS) and round-trip time (RTT) measurements. In the considered distance hijacking attack, a malicious UE communicates with a compromised benign UE over an ad-hoc WiFi network using CSMA, while the benign UE maintains a legitimate 5G NR connection with the gNodeB. The benign UE is configured to relay all messages between the malicious UE and the gNodeB without delay. The proximity threshold is set to 50 meters, in accordance with FCC guidelines [FCC_Thresholds], such that devices within this range are classified as proximate, while those beyond it are considered distant.
@@ 6 Related Work
(i) Location Privacy and Anonymity in DB-CRN: In DB-CRNs, regulatory requirements mandate that unlicensed users (i.e., SUs) disclose precise location information, device characteristics, and spectrum requests to geolocation databases in order to obtain authorized channels and transmission parameters [agarwal2022survey]. This regulatory enforcement fundamentally conflicts with the principle of privacy, as users must sacrifice their privacy for spectrum access, thereby exposing mobility patterns and long-term behavioral profiles, which then enables tracking and profiling attacks [grissa2021anonymous]. Moreover, location privacy without anonymity or unlinkability remains insufficient, as stable identifiers combined with coarse location allow tracking and linkage across queries [newman2022spectrum]. Standard protections (e.g., TLS) do not prevent legitimate entities from observing, storing, and correlating sensitive query metadata [shi2021challenges], either. Furthermore, the loss of anonymity expands the attack surface and can facilitate active threats such as selective DoS, targeted blocking, and user discrimination [weifang2010denial, darzi2024privacy].
@@
Despite extensive research efforts, existing privacy-enhancing mechanisms remain inadequate for real-world DB-CRN deployments. Schemes relying on k-anonymity or pseudonymous identifiers [zhu2019lightweight] provide only weak protection and lack rigorous security guarantees unless very large anonymity sets are maintained—an assumption that rarely holds in dense, mobile, or highly dynamic environments. Techniques based on location obfuscation or differential privacy [ul2022differential], while theoretically grounded, intentionally perturb location data and consequently impair the precision and dependability of spectrum availability decisions. Cryptographic approaches leveraging Private Information Retrieval (PIR) [grissa2021anonymous, xin2016privacy] offer stronger privacy guarantees but incur prohibitive computational and communication overhead, hindering scalability and practical deployment, while typically assuming benign user behavior and neglecting adversarial users. Consequently, our proposed solution utilizes primitives such as DAC and RLRS to provide effective protection in DB-CRNs through lightweight and scalable mechanisms that jointly ensure location privacy, anonymity, and unlinkability, while remaining compatible with regulatory mandates and operating in resource-constrained environments.
@@
(ii) Location Proof and Spoofing Resistance in DB-CRN: DB-CRNs fundamentally operate as location-based systems in which spectrum availability decisions depend on real-time, user-provided location information stored in geolocation DBs. Consequently, spoofed or falsified locations undermine the basis of regulatory authorization and can cause disproportionate harm, as attackers may impersonate legitimate entities or manipulate location and usage data to obtain unauthorized spectrum access, resulting in spectrum interference, service disruption, and economic loss [nguyen2019spoofing].
@@
While some approaches introduce location-proof mechanisms, they often rely on strong trust assumptions such as trusted infrastructure, specialized hardware, or dedicated location-verification servers, which are impractical in rural, ad hoc, or infrastructure-limited deployments [xin2016privacy]. The need for location verification becomes even more critical when combined with location privacy and anonymity, creating a fundamental paradox in DB-CRNs: revealing precise location information to prove correctness undermines privacy, while hiding location without verification enables spoofing. Thus, many existing solutions compromise user privacy by exposing fine-grained location data to verifiers, enabling tracking and linkage. These shortcomings highlight the necessity for a practical, privacy-preserving location verification framework that can authenticate user proximity without disclosing sensitive location data, while remaining robust against diverse spoofing and distance-fraud attacks under realistic DB-CRN deployment conditions [darzi2025slap].
@@
(iii) DoS Countermeasures and Spectrum Management for Next-Gen Network Systems: The centralized geolocation databases in DB-CRNs, due to their frequent interaction with massive IoT deployments, result in high-value choke points whose unavailability can disrupt spectrum access at scale [jakimoski2008denial]. Due to the cost asymmetry, even weak attackers can overwhelm spectrum DBs with excessive or illegitimate queries, causing high verification and processing overhead for servers, which in turn degrades spectrum coordination and reduces service availability.
Although various countermeasures have been proposed, including intrusion detection systems (IDSs), blockchain-based access control, and cryptographic Client Puzzle Protocols (CPPs), most existing defenses remain reactive and focus primarily on attack detection rather than prevention [darzi2024counter]. Such approaches often require continuous monitoring of traffic patterns, which incurs substantial operational overhead, and remain vulnerable to adversarial evasion [doriguzzi2024flad]. CPP-based defenses impose computational costs on clients prior to service access [bostanov2021client], thereby limiting the rate of malicious requests. However, at large scale, they introduce new bottlenecks related to puzzle generation, distribution, and parallelization, which can burden both servers and legitimate users [ali2020foundations]. While outsourcing puzzle generation can alleviate server-side overhead [darzi2024privacy], it merely shifts the attack surface to spectrum databases, which remain vulnerable to query flooding.
@@ A.1 Security Proofs
(i) Client Anonymity and Untraceability: SLAPX ensures client anonymity and untraceability via credential unforgeability and credential anonymity (Definitions 5 and 6). For any PPT adversary $\mathcal{A}$ interacting with issuance, delegation, and presentation oracles, the advantage in distinguishing which of two honest users generated a valid presentation satisfies $Adv_{\mathcal{A}}^{\mathrm{anon}}(\lambda):=\left|\Pr[b^{\prime}=b]-\tfrac{1}{2}\right|\leq negl(\lambda)$, where $b$ is the challenger’s hidden bit selecting the presenting user. Otherwise, $\mathcal{A}$ can be reduced to either (i) breaking the knowledge soundness of the underlying ZKPoK, (ii) distinguishing re-randomized structure-preserving signatures, or (iii) distinguishing public-key–switched presentations, yielding a distinguisher for the Decisional Diffie–Hellman (DDH) problem in the underlying group $\mathbb{G}$. In particular, the SPSEQ-UC scheme guarantees origin-hiding, derivation-privacy, and conversion-privacy [fuchsbauer2019structure, mir2023practical], implying that randomized signatures, delegated commitment vectors, and key-switched credentials are computationally indistinguishable from fresh ones. These properties compose under repeated application, ensuring that credential presentations remain anonymous and unlinkable with all but negligible probability.
(ii) Client Location Privacy: SLAPX preserves user location privacy through credential unlinkability (Definition 7). In particular, location information is protected by the unlinkability of signature–commitment pairs generated via the SPSEQ-UC scheme [mir2023practical]. The scheme supports signature re-randomization and public-key switching, ensuring that multiple presentations of the same underlying commitment yield computationally independent transcripts. Formally, for any PPT adversary $\mathcal{A}$ interacting with the proving oracle and observing a sequence of credential showings associated with the same location attribute $\ell$, the advantage of linking two presentations satisfies $Adv_{\mathcal{A}}^{\mathrm{link}}(\lambda):=\Big|\Pr[b^{\prime}=b]-\tfrac{1}{2}\Big|\leq negl(\lambda)$, where $b$ denotes the challenge bit selecting which of two unlinkable presentations is shown. Otherwise, $\mathcal{A}$ can be used to distinguish re-randomized SPSEQ-UC signatures or correlate switched public keys, contradicting the unlinkability of SPSEQ-UC under the generic group model. Consequently, repeated spectrum queries and location proofs in SLAPX remain computationally indistinguishable and unlinkable to verifiers, ensuring location privacy with all but negligible probability.